A military jet flies directly over and at right angles to the straight-line course of a commercial jet. The military jet is flying 200 mph faster than four times the commercial jet. How fast is each going if they are 2050 miles apart (on a direct line) after 1 hour?
Please Help! I have to show my work and I keep getting lost.

Answer by richwmiller(17219)

what was the point of the right angle?
Setting the stage for the pythagorean theorem.
the distance between them will be the hypotenuse
their flight patterns are the legs of the triangle.
(4x+200)^2+x^2=2050^2
17x^2+1600x+40000 = 4202500
(x-450)(17x+9250) = 0
we can only use the positive solution
x=450 commercial
military=4(450)+200=2000